OpenStudy (anonymous):

How do you solve 7.13((100-x)/100) I know I have to factor but its been a long time and don't remember how

OpenStudy (anonymous):

ok so first Multiply 7.13 by the terms inside the parentheses.

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

No thats not right. I have to factor the numerator. Parenthesis first.

OpenStudy (anonymous):

I can get to ((10-x)(10+x))/100 but don't know what to do with this part. The final step is to multiple by 7.13

OpenStudy (anonymous):

ok first can i show you my way then, you could decide if you wish to proceed and stick with it, but if not i could help you with urs? what do you say?

OpenStudy (anonymous):

ok. show me your way

OpenStudy (anonymous):

ok do you know that answer at all?

OpenStudy (anonymous):

nope. I am trying to solve for the percent of copper in a penny. The full equation is this: 6.23=8.96(x/100)+7.13((100-x)/100)

OpenStudy (anonymous):

Don't like it. That is completely wrong. Thanks anyway

OpenStudy (anonymous):

ok then show me your way :)

OpenStudy (tkhunny):

6.23=8.96(x/100)+7.13((100-x)/100) Multiply everything by 100 623 = 8.96x + 7.13(100-x) Distributive Property with the remaining parentheses 623 = 8.96x + 71.3 - 7.13x Subtract 71.3 551.7 = 8.96x - 7.13x Combine like terms on the right 551.7 = 1.83x Divide by 1.83 301.5 = x Well, if that's a percentage, something must be wrong with the setup. You may be confused by formulas including percentages. They can be very odd-looking and odd-working. If we adjust for that, we might get: 6.23=8.96x+7.13(100-x) which leads to yet another really bad result. Anyway, perhaps we could see exactly what it is you are doing?

OpenStudy (anonymous):

@tkhunny , i helped her and showed her how to solve 7.13((100-x)/100), since that is what she was asking for, and she did not like it, so i asked her how she would approach it, she never replied, and told me mine was wrong?? but her question is not clear of what she needs to solve for or do?

OpenStudy (tkhunny):

@some_someone Yes, I see that. It was a valiant effort. I was bored and thought it worth a little more effort, just to see if we could figure out the real problem, rather than work on just the asked problem. No offense intended. I can be a little annoying, at times.

OpenStudy (anonymous):

its okaii :)

OpenStudy (anonymous):

I got it figured out. Thanks anyways.